OpenStudy (anonymous):

A computer company features a new model with a larger 20-inch flat-panel display. The size is the diagonal length of the screen. Older models include 15 and 17 inch screens. 1) What are possible dimensions of the older 15-inch model? 2) Is the 20 inch model twice as large as the 15 inch one? 3) What model woud be twice as large as the 17 inch model?

OpenStudy (anonymous):

Looks like a Pythagorean question

OpenStudy (anonymous):

I think so, but they only give the diagonal.

OpenStudy (anonymous):



OpenStudy (anonymous):

Probably why I guess the question asks for possible dimensions

OpenStudy (anonymous):

Yeah, but how can I find the dimensions with onnly the diagonal?

OpenStudy (anonymous):

Well this means for problem 1: \[15^2 = a^2 + b^2\]

OpenStudy (anonymous):

I guess we can choose any dimensions as long as it equals this

OpenStudy (anonymous):

I guess so, but any?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

So I could do 10.6 for both?

OpenStudy (anonymous):

How did you get 10.6?

OpenStudy (anonymous):

15^ is 200 200 dvided by is 10.6

OpenStudy (anonymous):

Okay I see now

OpenStudy (anonymous):

That's the only way I can think of doing this

OpenStudy (anonymous):

@DanJS @triciaal @bibby

OpenStudy (anonymous):

I think if we assume that the other sides are equivalent is the only way

OpenStudy (anonymous):

So if we do this then we can just do this for the sides now: \[c^2 = 2a^2\]

OpenStudy (anonymous):

So the only answer I can think for number 1 is a 10.6 in. by 10.6 in. screen

OpenStudy (triciaal):

if you meant 15^2 then that would be 225 not 200

OpenStudy (anonymous):

I get 10.6 with 225 so it seems like a typo for the same number in mind

OpenStudy (triciaal):

i was thinking of a ratio the new is 20 the old is 15 4:3

OpenStudy (triciaal):